2.A balloon containing 7.83 moles of He has a volume of 27.5 L at a certain temperature and pressure. What would be the volume of the balloon, in liters, if 1.00 mole of Heis removed from the balloon at constant temperature and pressure?
Answer #1

Given:

Vi = 27.5 L

ni = 7.83 mol

nf = 6.83 mol

use:

Vi/ni = Vf/nf

27.5 L / 7.83 mol = Vf / 6.83 mol

Vf = 24.0 L

Answer: 24.0 L